OpenStudy (anonymous):

No2 (nitrogen dioxide) and N2O4 (dinitrogen tetroxide) have the same empirical formula, NO2. Confirm this by calculating the percent by mass of each element present in the two compounds.

OpenStudy (anonymous):

N2O4= 92 g. N2= 28g. O4=64 g. 28/92 x100=30.4% 64/92x100=70% ---------------------------------------- NO2= 46g. N=14 g. O=32 g. 14/46x100=30.4% 32/46x100=70%

OpenStudy (anonymous):

Let me know if you still need help :)

OpenStudy (anonymous):

Thank you!